Velocity

The velocity graph of a car accelerating from rest to a speed of 60 km/h over a period of 30 seconds is shown. Estimate the distance traveled during this period. (UseM6 to get the most precise estimate. Round your answer to two decimal places.)

km=
Answer #1
average velocity is ( Vf + 0 )/2 = 75/2 = 37/5 kph

Distance is V x t = but 30 is in seconds.

30 seconds x 1 min/60 seconds x 1 hr/60 min = 0.0083 hr

Distance = 37.5 x 0.0083 = .312 km < ========= Ans
answered by: please please please helpppp
Answer #2

0.36 is the correct answer

source: t
answered by: kay